OpenStudy (anonymous):

a bag of pins has 4 black, 2 brown and 4 blue pins in it. if 3 pins are selected without replacement then determine the probability of selecting exactly 1 brown? pls help :)

OpenStudy (amriju):

There can be three cases...1 brown with...i) two black...ii)two blue..and iii) one black n one blue..... probabbility of just one brown=2/10....now probbability that (i) occurs..is probb of 1 brown and prob of two black thats 4C2/10C2=2/15...so net prob for (i)=2/15*1/5=2/75...similar for case (ii)...case (iii) is 4/10*4/10*2/10...now all these cases r independent..add them..and thats the probbability

OpenStudy (anonymous):

thanak you for your help